# A Random Rule Model

### 04 Mar 2026 v1

@@ 9.3 Cross-fitted library restrictions
For each CV fold, we fit the full rule-gating model (all 12 rules, including the attention rules) on the training set and compute the average effective responsibility $w_{f}$ on the training data. We then rank all 12 rules by $w_{f}$ and select the top-$k$ (for $k\in\{3,5,7,12\}$). For each $k$, we refit a restricted gate using only the selected rules on the same training set and evaluate on the held-out test set. The selection is “cross-fitted” in the sense that rule subsets are chosen using in-sample information only, and the restricted model’s performance is evaluated strictly out of sample.
@@
An alternative selection unit is the behavioral family: we group rules into five predefined families - extremum (MMn, MMa, MMx, MAP), salience (SAL, SAL2), regret (REG, REGmed), disappointment (DIS, DISmed), and attention (A1, A2) - and select the top-$k_{\mathrm{fam}}$ families (for $k_{\mathrm{fam}}\in\{1,2,3,4,5\}$) by aggregate family responsibility on the training set. All rules within the selected families are included, and performance is again evaluated out of sample.
Table 5 reports results. The main finding is that performance retention is high even with substantial library reduction. Under top-$k$ selection, restricting to $k=7$ rules retains the bulk of the full-library predictive accuracy, with $k=5$ still substantially outperforming a constant predictor, although at this level the model no longer clearly dominates all structured neural benchmarks. At $k=3$, retention drops to about $25\%$, indicating that three rules are too few to span the behavioral variation in the data. Under family selection, two families suffice for strong performance, but restricting to a single family produces negative retention (MSE exceeding twice the full-library level), because no single family can substitute for the complementary mechanisms provided by the others. Selection is stable across folds: the rules that the full model weights most heavily are also the ones most frequently selected out of sample (see Appendix Table A.6 for fold-by-fold selection frequencies).
These results address the “handpicked library” concern in two ways. First, the identity of the selected rules is consistent with the main mechanism results in Section 8: the top-ranked rules in cross-fitted selection match the highest-responsibility rules in the full model. Second, the fact that small subsets retain strong performance confirms that the library’s predictive contribution is concentrated on a few core mechanisms rather than distributed thinly across many rules.
@@
Full (12 rules) 12 0.01248 0.00008 100.0
Top-$k$ 7 0.01381 0.00010 89.3
Top-$k$ 5 0.01548 0.00013 75.9
Top-$k$ 3 0.02185 0.00020 24.9
Full (5 families) 5 0.01247 0.00007 100.0
Top-family 4 0.01268 0.00008 98.3
Top-family 3 0.01368 0.00010 90.3
Top-family 2 0.01596 0.00012 72.0
Top-family 1 0.02736 0.00024 -19.4
@@ 9.4 Attention-rule removal
The attention rules absorb about one-third of effective responsibility ($w_{\textsc{A1}}\approx 0.193$, $w_{\textsc{A2}}\approx 0.146$), capturing a channel that is plausibly important in MTurk data where inattention and default-like behavior are prevalent. Table A.1 (Appendix) replicates the main cross-validation with the 10-rule model (excluding A1/A2). The 10-rule model achieves MSE $0.0201$, outperforming Neural EU ($0.0222$), Neural CPT ($0.0215$), and Neural PT ($0.0206$), while remaining close to the value-based network ($0.0192$). This confirms that the attention channel carries substantial predictive content (MSE rises from $0.0128$ to $0.0201$), but the non-attention rules alone still surpass most structured benchmarks.
The qualitative context-dependent patterns also persist under the 10-rule specification: MMx rises with complexity while MAP declines, and along risk asymmetry, extreme-outcome comparisons decrease while MAP increases; see Appendix C, Figures A.1A.2.
@@ 9.5 Varying the activity discipline
Our baseline activity discipline requires strict first-order stochastic dominance (FOSD) of the conduct-transformed lotteries. This threshold is sharp: a rule is active whenever the survival function of the recommended option weakly dominates and strictly exceeds the other at some point. We probe the sensitivity of the model to this discipline in two directions.
Strengthening the discipline ($\varepsilon>0$).
We require an $\varepsilon$-gap in the survival functions: specifically, $S^{1}(z)\geq S^{2}(z)+\varepsilon$ for all grid points $z$ (beyond the common starting point) and $S^{1}(z)>S^{2}(z)+\varepsilon$ for at least one grid point.
For $\varepsilon\in\{0.01,0.05,0.10\}$, we rebuild the rule activity indicators and re-run the full 50-fold cross-validation protocol. Table 6 reports, for each $\varepsilon$, the mean number of active rules per menu, the average overlap with the majority choice direction, and the out-of-sample MSE from the same CV protocol as Table 1.
As $\varepsilon$ increases, the activity discipline becomes more stringent: fewer rules are active per menu, and the remaining active rules tend to be those with clearer dominance relationships. The key finding is that the overall structure of the activity pattern - in particular, the relative ranking of rules by mean activity - is robust to moderate increases in $\varepsilon$, although the absolute level of activity naturally decreases. On the predictive side, the MSE at $\varepsilon=0.01$ remains close to the baseline, confirming that the model’s predictions do not hinge on marginal dominance relationships. At larger $\varepsilon$, fewer rules are available on each menu, and MSE increases, reflecting the loss of discriminatory power as the library becomes more conservative.
Removing the discipline ($\varepsilon<0$).
A complementary question is whether the FOSD-based activation itself carries predictive information, or whether the gate could achieve similar performance if all rules were declared active on every menu. To test this, we set $\varepsilon=-2$, which makes the dominance condition trivially satisfied: all 12 rules are active on every menu (mean active rules =12=12). Recommendations for rules that have standard FOSD are preserved; rules for which neither direction dominates are assigned a default (right-option) recommendation, adding noise. All other aspects of the model - architecture, learning rate, training horizon - are held fixed.
Table 6 includes this specification. Without the activity discipline, MSE increases only modestly relative to the baseline ($0.01294$ vs. $0.01282$), indicating that the gate is flexible enough to partially compensate by down-weighting rules that lack a clear dominance signal. Notably, the no-discipline specification ($\varepsilon=-2$) still outperforms all positive-$\varepsilon$ variants, where genuinely informative rules are progressively removed. The pattern across the full range of $\varepsilon$ values is therefore asymmetric: over-restricting the activity set (high $\varepsilon$) is more costly than under-restricting it (all rules active), because the former removes rules that carry genuine predictive content, while the latter only adds noise that the gate can learn to suppress.
This asymmetry clarifies the role of the FOSD-based activation: for prediction, the gate can largely learn on its own which rules to suppress, but for interpretation, the activation criterion ensures that responsibility weights are nonzero only on menus where the rule delivers a decisive ranking. The activation discipline is therefore best understood as an interpretive discipline rather than a predictive one.
@@ 9.6 Placebo rule library, feature-based baselines, and raw-encoding gate
Three additional robustness checks are reported in Appendix LN. First, a placebo test scrambles each rule’s indicators across menus within complexity strata, preserving marginal activity and recommendation rates but destroying the menu-specific mapping. The placebo library produces substantially higher MSE than the real library (Appendix Table A.8), confirming that the rule indicators carry genuine, menu-specific predictive content.
@@
Third, a raw-encoding gate variant uses the full $40$-dimensional menu representation $\psi(A)$ instead of $z(A)$ (Appendix Figure A.5). It achieves MSE $0.01464$, about $14\%$ worse than the baseline $z(A)$ specification ($0.01282$), confirming that the interpretable summary statistics capture decision-relevant variation without sacrificing accuracy. The global identification conditions (G1)–(G2) are not applicable to this parameterization.
Summary of robustness evidence
Taken together, these checks confirm that the gate is globally identified, that the decomposition is stable across training samples, and that the rule indicators carry genuine predictive content. Performance is not an artifact of the ex ante library choice, the attention channel, or the specific activity discipline. The interpretable gate features $z(A)$ capture decision-relevant variation at least as well as the full raw encoding. These results support interpreting the RRM’s decompositions as disciplined, data-supported statements about procedural variation.
@@ 10 Conclusion
This paper develops a procedural approach to representing how behavior varies across a rich environment space. We introduce a Random Rule Model (RRM), in which predicted behavior is generated by switching among a small library of transparent, deterministic decision rules, disciplined by an activity criterion. A differentiable gating mechanism maps interpretable summary statistics into rule propensities, producing an explicitly interpretable mixture over named lines of conduct. We develop an identification theory for RRMs, showing that global identification follows from two verifiable conditions on the observed menu support - within-feature variation in decisive-side patterns and affine richness - both directly checkable on the data.
Applied to 9,8319{,}831 binary lottery menus from the choices13k dataset, rule-gating substantially outperforms structured neural benchmarks, closing about $87\%$ of the baseline-to-MOT (Mixture of Theories) gap while remaining highly restrictive under permutation-fit tests. The model exports well to CPC18 without re-tuning, in contrast to the most flexible benchmark, which deteriorates sharply out of domain. Mechanism diagnostics reveal that a small subset of procedures - extreme-outcome screening, salience-based comparisons, and modal-outcome comparisons - drives most of the predictive content, with systematic reallocation along complexity and risk asymmetry. Both identification conditions are verified on the data, and a battery of robustness checks confirms that findings are not driven by the ex-ante library choice, marginal dominance relationships, or generic model flexibility.
More broadly, the RRM brings economic interpretability into high-dimensional prediction without reverting to a single valuation model. The framework may be portable beyond risky choice to settings with rich environment spaces and repeated behavior: consumer choice, dynamic decision problems, strategic settings, or social choice. Natural extensions include enlarging rule libraries, incorporating richer heterogeneity, and extending the identification conditions to richer gate parameterizations.
@@ A.1 Choice environment
Assumption 1 (Interior probabilities).
@@ A.6 Identification map and sufficient-variation condition
For each multi-sided menu $A$, fix a reference alternative $i_{0}(A)\in A$ (any alternative with $\mathcal{I}_{i_{0}}(A)\neq\emptyset$). By Assumption 1 and (A.5), the log-ratios
@@
are well defined for each $i\in A\setminus\{i_{0}(A)\}$.
@@ A.7 Global identification theorem
Theorem 2 (Global identification of the general CA-RRM).
Consider the CA-RRM with affine indices $u_{f}(A)=\alpha_{f}+\beta_{f}^{\top}\psi(A)$, $f\in\mathcal{F}$. Assume $P(i\mid A)\in(0,1)$ for all $i\in A$ and all menus in the support (Assumption 1), and impose the normalization (A.6). Let $d=\dim(\psi)$. Suppose there exist $(d+1)$ feature values $x^{(0)},\ldots,x^{(d)}$ such that:
@@
The proof parallels that of Theorem 1 in the main text, extended to the multinomial setting.
@@
Remark 1.
The binary application in the main text (Theorem 1) is the special case with $|A|=2$, where each menu produces a single row $h(A)^{\top}$ and the odds ratio reduces to p(A)/(1p(A))p(A)/(1-p(A)).
@@ A.8 Local identification theorem
Remark 2.
Theorem 2 above establishes global identification under conditions (G1)–(G2), which require that multiple menus share the same feature value. The following local result provides an alternative characterization based on the Jacobian of the identification map. It is particularly relevant when the gate operates on a high-dimensional feature map (such as the raw menu encoding $\psi(A)\in\mathbb{R}^{4M}$) for which the replication condition in (G1) may not hold. See Appendix O for further discussion.
Theorem 3 (Local identification of the general CA-RRM).
@@
For part (a): if $J(\vartheta^{\star})$ has rank $K$, then by the inverse function theorem (applied to any $K$-dimensional subvector of $H$ whose sub-Jacobian has rank $K$), $H$ is locally injective at $\vartheta^{\star}$. Hence $H(\vartheta)=H(\vartheta^{\star})$ implies $\vartheta=\vartheta^{\star}$ in a neighborhood.
@@ A.9 Closed-form Jacobian
We derive the Jacobian entries in closed form. For each active rule $f\in\mathcal{F}_{A}$, define the within-side softmax share
@@
Proposition 4 (Jacobian entries).
@@
which recovers the Jacobian entries of the main text. The binary global identification theorem in the main text (Theorem 1) is the special case of Theorem 2 obtained by setting $|A|=2$.
@@
Appendix B Proof of Theorem 1

### 19 Mar 2026 v2

@@ 8.1 Cross-fitted library restrictions
For each CV fold, we fit the full 12-rule model on the training set, rank rules by $w_{f}$, select the top-$k$ ($k\in\{3,5,7,12\}$), refit a restricted gate, and evaluate out of sample.
@@
We group rules into five families—extremum, salience, regret, disappointment, attention—and select the top-$k_{\mathrm{fam}}$ families ($k_{\mathrm{fam}}\in\{1,\ldots,5\}$) by aggregate training-set responsibility.
Table 6 reports results. Restricting to $k=7$ rules retains the bulk of full-library accuracy; at $k=3$, retention drops to about $34\%$, indicating that three rules are too few to span the behavioral variation. Under family selection, two families suffice for strong performance, but a single family does not. Selection is stable across folds (Appendix Table A.6), and the selected rules match the highest-responsibility rules in the full model, confirming that the library’s predictive content is concentrated on a few core mechanisms.
@@
Full (12 rules) 12 0.01199 0.00007 100.0
Top-$k$ 7 0.01266 0.00008 94.4
Top-$k$ 5 0.01388 0.00010 84.2
Top-$k$ 3 0.01992 0.00013 33.9
Full (5 families) 5 0.01198 0.00007 100.0
Top-family 4 0.01214 0.00007 98.6
Top-family 3 0.01324 0.00009 89.5
Top-family 2 0.01484 0.00009 76.1
Top-family 1 0.02304 0.00013 7.6
@@ 8.2 Attention-rule removal
As documented in Section 7.3, the attention rules absorb roughly half of effective responsibility, making them an important channel to scrutinize. Table A.1 (Appendix) replicates the main cross-validation with the 10-rule model (excluding A1/A2). The 10-rule model achieves MSE $0.0201$, outperforming Neural EU ($0.0222$), Neural CPT ($0.0215$), and Neural PT ($0.0206$), while remaining close to the value-based network ($0.0192$). This confirms that the attention channel carries substantial predictive content (MSE rises from $0.0117$ to $0.0201$), but the non-attention rules alone still surpass most structured benchmarks. The context-dependent patterns documented in Figures 56 also persist qualitatively under the 10-rule specification; see Appendix E.
@@ 8.3 Varying the activity discipline
Our baseline requires strict FSD of the conduct-transformed lotteries. We probe sensitivity by requiring an $\varepsilon$-gap in survival functions for $\varepsilon\in\{0.01,0.05,0.10\}$ (strengthening the discipline) and by setting $\varepsilon=-2$ (all 12 rules active on every menu, removing the discipline entirely).
Table 7 reports results. As $\varepsilon$ increases, fewer rules are active per menu and MSE rises, reflecting lost discriminatory power. At $\varepsilon=0.01$, MSE remains close to the baseline, confirming that predictions do not hinge on marginal dominance. Without the discipline ($\varepsilon=-2$), MSE increases only modestly ($0.01202$ vs. $0.01168$): the gate compensates by down-weighting rules lacking a clear signal. The pattern is asymmetric: over-restricting the activity set (high $\varepsilon$) is more costly than under-restricting it, because the former removes genuinely informative rules while the latter only adds noise the gate can suppress. The FSD-based activation is therefore best understood as an interpretive discipline ensuring responsibility weights are nonzero only where a rule delivers a decisive ranking rather than a predictive one.
@@ 8.4 Placebo rule library, feature-based baselines, and raw-encoding gate
Three additional robustness checks are reported in Appendix MN. First, a placebo test scrambles each rule’s indicators across menus within complexity strata, preserving marginal activity and recommendation rates but destroying the menu-specific mapping. The placebo library produces substantially higher MSE than the real library (Appendix Table A.8), confirming that the rule indicators carry genuine, menu-specific predictive content.
@@
Third, a raw-encoding gate variant uses the full $40$-dimensional menu representation $\psi(A)$ instead of $z(A)$ (Appendix Figure A.5). It achieves MSE $0.0139$, about $19\%$ worse than the baseline $z(A)$ specification ($0.0117$), confirming that the interpretable summary statistics capture decision-relevant variation without sacrificing accuracy. The global identification conditions (G1)–(G2) are not applicable to this parameterization.
Taken together, these checks confirm that the decomposition is stable, the rule indicators carry genuine predictive content, and the results are not artifacts of the library choice, attention channel, activity discipline, or gate features.
@@ 9 Conclusion
This paper studies stochastic choice under latent procedural heterogeneity. We develop a Random Rule Model (RRM) in which observed choice probabilities arise from switching among a small library of transparent deterministic decision procedures, with environment-dependent rule weights. The main contribution is an identification-and-estimation framework for recovering these latent procedural weights from menu-level choice frequencies. Identification has a two-step structure: within-feature variation in decisive-side patterns identifies relative rule weights, while cross-feature affine richness identifies the gate parameters. A constructive two-step estimator then supports inference for economically meaningful rule-level functionals.
Applied to risky choice, the framework uncovers substantial and systematic procedural heterogeneity. Salience-based comparisons, regret-type rules, and modal-outcome comparisons receive the largest responsibility weights among structured procedures, and rule responsibility shifts predictably with menu characteristics such as tradeoff complexity and dispersion asymmetry. These comparative statics provide an empirical characterization of risky choice in terms of changing procedure use rather than only reduced-form choice frequencies.
The predictive exercises support the validity of the representation. Rule-gating substantially outperforms structured neural benchmarks, remains highly restrictive under permutation-fit diagnostics, and retains predictive content on an independent dataset. At the same time, the evidence from varying the activity discipline shows that the activity filter is best understood as an attribution operator that restricts behavioral responsibility to rules that are actually decisive, rather than as the primary source of fit.
On the econometric side, the two-step overidentification test rejects the affine gate specification for most rules, indicating that the parsimonious linear-index parameterization does not fully capture how rule weights vary across menu characteristics. Since the first-stage identification is nonparametric, a natural extension is to replace the affine second step with a richer specification. More broadly, the paper suggests that stochastic choice data can be used to recover not only how often each option is chosen, but also how populations reallocate across decision procedures as environments change. This makes procedural heterogeneity a tractable econometric object, and opens the door to applications in consumer choice, dynamic choice, and strategic environments where behavior may be governed by shifting mixtures of transparent rules.
@@ A.1 Choice environment
Assumption A.1 (Interior probabilities).
@@ A.6 Identification map and sufficient-variation condition
For each multi-sided menu $A$, restrict attention to alternatives $i\in A$ with $\mathcal{I}_{i}(A)\neq\emptyset$ (i.e., alternatives recommended by at least one active rule). [^13] Under correct specification, Assumption A.1 implies $P(i\mid A)>0$ for all $i\in A$, which requires $\mathcal{I}_{i}(A)\neq\emptyset$ since (A.5) assigns zero probability to uncovered alternatives. Hence the restriction to covered alternatives is implicit in the joint assumption of correct specification and interior probabilities. Fix a reference alternative $i_{0}(A)$ among these. By Assumption A.1 and (A.5), the log-ratios
@@
are well defined for each covered $i\in A\setminus\{i_{0}(A)\}$.
@@ A.7 Global identification theorem
Theorem A.1 (Global identification of the general CA-RRM).
Consider the CA-RRM with affine indices $u_{f}(A)=\alpha_{f}+\beta_{f}^{\top}\psi(A)$, $f\in\mathcal{F}$. Assume $P(i\mid A)\in(0,1)$ for all $i\in A$ and all menus in the support (Assumption A.1), and impose the normalization (A.6). Note that under correct specification, this interior-probability condition implies that every alternative $i\in A$ is recommended by at least one active rule (i.e., $\mathcal{I}_{i}(A)\neq\emptyset$), since (A.5) assigns zero probability to uncovered alternatives. Let $d=\dim(\psi)$. Suppose there exist $(d+1)$ feature values $x^{(0)},\ldots,x^{(d)}$ such that:
@@
The proof parallels that of Theorem 1 and Corollary 2 in the main text, extended to the multinomial setting.
@@
Remark A.1.
The binary application in the main text (Corollary 2) is the special case with $|A|=2$, where each menu produces a single row $h(A)^{\top}$ and the odds ratio reduces to p(A)/(1p(A))p(A)/(1-p(A)).
@@ A.8 Local identification theorem
Remark A.2.
Theorem A.1 above establishes global identification under conditions (G1)–(G2), which require that multiple menus share the same feature value. The following local result provides an alternative characterization based on the Jacobian of the identification map. It is particularly relevant when the gate operates on a high-dimensional feature map (such as the raw menu encoding $\psi(A)\in\mathbb{R}^{4M}$) for which the replication condition in (G1) may not hold. See Appendix O for further discussion.
Theorem A.2 (Local identification of the general CA-RRM).
@@
For part (a): if $J(\vartheta^{\star})\in\mathbb{R}^{D\times K}$ has rank $K$, there exist $K$ components of $H$ whose $K\times K$ sub-Jacobian is nonsingular (select any $K$ linearly independent rows of $J(\vartheta^{\star})$). By the inverse function theorem applied to this square submap, the selected components of $H$ are locally injective at $\vartheta^{\star}$. Since $H(\vartheta)=H(\vartheta^{\star})$ implies equality of the selected subvector, the full map $H$ is locally injective.
@@ A.9 Closed-form Jacobian
We derive the Jacobian entries in closed form. For each active rule $f\in\mathcal{F}_{A}$, define the active-set softmax share (i.e., the share of rule $f$ in the softmax over all active rules, not conditional on alternative)
@@
Proposition A.3 (Jacobian entries).
@@
which is the binary specialization of the general Jacobian entries in Proposition A.3. The binary global identification result (Corollary 2) is the special case of Theorem A.1 obtained by setting $|A|=2$.
@@
Appendix B Proof of Theorem 1 (fixed-feature identification)
Proof.
Fix $x\in\mathbb{R}^{d}$ and let $\mathcal{A}_{2}(x):=\{A\in\mathcal{A}_{2}:\psi(A)=x\}$. For each menu $A\in\mathcal{A}_{2}(x)$, define the odds ratio r(A):=p(A)/(1p(A))r(A):=p(A)/(1-p(A)) and the column vector $h(A)\in\mathbb{R}^{|\mathcal{F}|}$ with entries $h_{f}(A)=\kappa_{f}^{L}(A)-r(A)\kappa_{f}^{R}(A)$. Denote the true weights by $\omega_{f}(x)=\exp(\alpha_{f}+\beta_{f}^{\top}x)>0$.
Step 1 (from choice probabilities to linear restrictions). Under the CA-RRM the model-implied left-choice probability at menu $A$ is
$p(A)=\frac{\sum_{f}\kappa_{f}^{L}(A)\,\omega_{f}(x)}{\sum_{f}\kappa_{f}^{L}(A)\,\omega_{f}(x)+\sum_{f}\kappa_{f}^{R}(A)\,\omega_{f}(x)}.$
Cross-multiplying $p(A)\sum_{f}\kappa_{f}^{R}(A)\omega_{f}(x)=(1-p(A))\sum_{f}\kappa_{f}^{L}(A)\omega_{f}(x)$ and dividing by (1p(A))(1-p(A)) gives
$\sum_{f\in\mathcal{F}}\bigl(\kappa_{f}^{L}(A)-r(A)\,\kappa_{f}^{R}(A)\bigr)\,\omega_{f}(x)\;=\;h(A)^{\top}\omega(x)\;=\;0.$
Since this holds for every $A\in\mathcal{A}_{2}(x)$, any vector in the cellwise identified set $\Omega(x)$ lies in the null space of the matrix $H(x)$ formed by stacking the rows $\{h(A)^{\top}:A\in\mathcal{A}_{2}(x)\}$.
Step 2 (sufficiency: $\operatorname{rank}H(x)=|\mathcal{F}|-1$ $\Rightarrow$ single ray). Suppose $\operatorname{rank}H(x)=|\mathcal{F}|-1$. Then $\ker H(x)$ is one-dimensional. Let $v$ be a nonzero vector spanning this kernel. Because the true weights $\omega(x)\in\mathbb{R}_{++}^{|\mathcal{F}|}$ lie in $\Omega(x)\subseteq\ker H(x)$, we have $\omega(x)=c\,v$ for some scalar $c\neq 0$. Since $\omega(x)$ is strictly positive, $v$ must be a scalar multiple of a strictly positive vector. Hence every $\tilde{\omega}\in\Omega(x)$ satisfies $\tilde{\omega}=c^{\prime}\,v$ for some $c^{\prime}>0$, i.e. $\Omega(x)=\{c^{\prime}\,\omega^{\star}(x):c^{\prime}>0\}$ is a single positive ray.
Step 3 (necessity: single ray $\Rightarrow$ $\operatorname{rank}H(x)=|\mathcal{F}|-1$). We prove the contrapositive. Suppose $\operatorname{rank}H(x)<|\mathcal{F}|-1$. Then $\dim\ker H(x)\geq 2$. Pick two linearly independent vectors $v_{1},v_{2}\in\ker H(x)$ with $v_{1}=\omega(x)\in\mathbb{R}_{++}^{|\mathcal{F}|}$. For all sufficiently small $\varepsilon>0$, the vector $v_{1}+\varepsilon\,v_{2}$ is also strictly positive (since $v_{1}$ has strictly positive entries). Moreover $v_{1}+\varepsilon v_{2}\in\ker H(x)$ by linearity. Therefore $v_{1}+\varepsilon v_{2}\in\Omega(x)$, and it is not proportional to $v_{1}$ (since $v_{1}$ and $v_{2}$ are linearly independent), so $\Omega(x)$ is not a single ray.
Step 4 (point identification under normalization). Normalizing $\omega_{f_{0}}(x)=1$ restricts $\Omega(x)$ to the hyperplane $\{\omega:\omega_{f_{0}}=1\}$. A single positive ray intersects this hyperplane in a unique point, so the normalization converts ray identification into point identification. ∎
Appendix C Proof of Corollary 2 (detailed version)